OpenStudy (el_arrow):

find the directional derivative of the function at the given point in the direction of the vector v

OpenStudy (el_arrow):

g(s,t) = s*sqrt(t) , P(2,4), v=2i-j

OpenStudy (el_arrow):

i got gs and gt. gs= sqrt(t) and gt= s/(2sqrt(t))

OpenStudy (el_arrow):

and u = <1/sqrt(5) , -1/(2sqrt(5))

OpenStudy (el_arrow):

the answer on the back of the book is 7/(2sqrt(5)) but im getting a different answer

OpenStudy (el_arrow):

@myininaya @zepdrix

OpenStudy (el_arrow):

@wio

OpenStudy (anonymous):

Do you know how to find the gradient?

OpenStudy (el_arrow):

hey man i keep getting a different answer than the one from the book could you help me out real quick

OpenStudy (el_arrow):

yeah i multiplied the u * gs +u*gt

OpenStudy (anonymous):

\[ g(s,t) = s\cdot \sqrt{t} \implies \nabla g = \sqrt{t}~\mathbf i+\frac{s}{\sqrt t}~\mathbf j \]

OpenStudy (anonymous):

\[ D_{\mathbf v}g = \nabla g\cdot \mathbf v= 2\sqrt t -\frac{s}{\sqrt t} \]

OpenStudy (el_arrow):

yeah its \[\frac{ s }{ 2\sqrt{t} }\]

OpenStudy (el_arrow):

for the j

OpenStudy (anonymous):

\[ \ldots = \frac{2t}{\sqrt t}-\frac{s}{\sqrt t} = \frac{2t-s}{\sqrt t} \]

OpenStudy (el_arrow):

okay

OpenStudy (anonymous):

Is that what you got for \(D_{\mathbf v}g(s,t)\)?

OpenStudy (el_arrow):

no i got 1/sqrt(5) * (sqrt(t)) -1/2sqrt(5)* (s/2sqrt(t))

OpenStudy (anonymous):

What is that even supposed to be?

OpenStudy (anonymous):

You should still have variables because you don't plug points in at this stage

OpenStudy (anonymous):

Did you find \(\nabla g(\mathbf u)\) rfirst?

OpenStudy (el_arrow):

yeah i got \[\frac{ 2 }{ 2\sqrt{5}} and -\frac{ 1}{ 2\sqrt{5} }\]

OpenStudy (el_arrow):

is it right?

OpenStudy (anonymous):

So you're saying \(t=5\) and \(s=1\)?

OpenStudy (anonymous):

I think your gradient is correct, but you're not plugging \(u\) in correctly, or you aren't telling me what the correct \(u\) is.